On Python 3.7.2:

How would I make functions for these string functions/methods without using the actual functions?

(For example, how would I find if a string is comprised of alphabets without using the function "isalpha()" ?)

len()

isalpha()

isupper()

isdigit()

swapcase()

string_lower()
Answer #1

OUTPUT :

CODE :

def len(x):
ans = 0
for c in x:
ans+=1
return ans

def isalpha(x):
for c in x:
if not ((c>='a' and c<='z') or (c>='A' and c<='Z')):
return False
return True

def islower(x):
for c in x:
if not (c>='a' and c<='z'):
return False
return True

def isupper(x):
for c in x:
if not (c>='A' and c<='Z'):
return False
return True

def isdigit(x):
for c in x:
if not (c>='0' and c<='9'):
return False
return True

def swapcase(x):
ans = ''
for c in x:
if isupper(c):
ans = ans + chr(ord(c)+32)
elif islower(c):
ans = ans + chr(ord(c)-32)
else:
ans = ans+c
return ans

def string_lower(x):
ans = ''
for c in x:
if isupper(c):
ans = ans + chr(ord(c)+32)
else:
ans = ans+c
return ans